# Coloring outside the lines: Spectral bounds for generalized hypergraph colorings

### v1

@@ 1 Introduction
To accommodate more flexible coloring constraints, De Werra (1979) [12] introduced $I$-regular colorings: for every edge, the number of vertices of any color that it contains is bounded both above and below. In this paper, we introduce and study a specific case of $I$-regular colorings called $d$-proper colorings: these are colorings where every edge contains at most $d$ vertices of a fixed color. The concept of $d$-proper colorings generalizes strong colorings (which correspond to the case $d=1$). Moreover, for $c$-uniform hypergraphs (i.e., hypergraphs whose edges all have the same size $c$), a $c-1$-proper coloring corresponds to a weak coloring.
@@
In this paper, we shall study spectral bounds for various notions of coloring numbers, both for classical hypergraphs and oriented hypergraphs, which were introduced by Shi (1992) [35], as a generalization of classical hypergraphs in which an orientation is attached to every vertex-edge incidence. The adjacency matrix and Kirchoff Laplacian matrices for oriented hypergraphs were introduced by Reff and Rusnak (2012) [30], and the normalized Laplacian was introduced by Jost and Mulas (2019) [21]. These matrices encode information about the structure of the hypergraphs, as they do in the graph case. We refer the reader to [7, 8, 13, 32, 18, 33, 22, 28, 29, 30, 34, 31] for literature on the adjacency and Kirchhoff Laplacian matrices, and to [3, 21, 24, 26, 25] for literature on the normalized Laplacian matrix.
@@ 2.1 The normalized Laplacian for hypergraphs
where the incidence matrix corresponds to any arbitrary orientation of the graph, such that each edge has exactly one input and one output. Equivalently, the graph Kirchhoff Laplacian can be written as $K=D-A$, where $D$ is the degree matrix, whose diagonal entries are the degrees of the corresponding vertices, and $A$ is the adjacency matrix. Similarly, the normalized Laplacian of a graph $G$ is the matrix
@@
[…]
@@
$L_{i,j}=\frac{\text{ edges where $v_{i}$ and $v_{j}$ are co-oriented }-\#\text{ edges where $v_{i}$ and $v_{j}$ are anti-oriented}}{\deg v_{i}}.$
@@
[…]
@@
Let $H$ be an $N$-dimensional vector space with a positive definite scalar product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ), and let $A\colon H\rightarrow H$ be a self-adjoint linear operator. Let $\mathcal{H}_{k}$ be the family of all $k$-dimensional subspaces of $H$. Then the eigenvalues $\lambda_{1}\leq\ldots\leq\lambda_{N}$ of $A$ can be obtained by
@@ 2.2 Coloring hypergraphs
In particular, if $\Gamma$ is a $c$-uniform hypergraph in which every vertex is an input in all edges it belongs to, then any $d$-proper coloring is also a $(d-1)$-tailored coloring. This is why $q$-tailored colorings generalize $d$-proper colorings in this setting.
@@ 3.1 Strong vertex coloring of general oriented hypergraphs
Our proof relies on Rayleigh quotients and leads to new consequences, which are detailed in Proposition 3.2 and 3.3. These results show that, if equality holds in the bound from Theorem 3.1, then there must a specific relation between the eigenspaces corresponding to $\lambda_{1}$ and $\lambda_{N}$. To better understand the structure of such hypergraphs, one needs to know these eigenspaces. This is why, to gain further insight, in Subsection 3.2, we shall focus on $c$-uniform unoriented hypergraphs, for which the eigenspace associated with $\lambda_{N}=c$ is known to consist of constant functions.
@@
=$\overset{(i)}{=}1+\bigl{(}l-1\bigr{)}\bigl{(}1-\textrm{RQ}(g_{12})\bigr{)}.$
@@
=$\overset{(ii)}{=}1-\bigl{(}k-1\bigr{)}\bigl{(}\textrm{RQ}(h_{12})-1\bigr{)}.$
@@
    For every eigenfunction $g$ with eigenvalue $\lambda_{N}$, and for all proper $\chi$-coloring with corresponding coloring classes $V_{1},\ldots,V_{\chi}$, we have
@@ 3.2 Strong vertex coloring of unoriented $c$-uniform hypergraphs
    If $\chi=c$, then $\lambda_{1}=0$, and the eigenvalue 00 has multiplicity at least $c-1$.
@@
Since $\lambda_{1}\geq 0$, the bound must be sharp. By Proposition 3.2, this implies that the multiplicity of the eigenvalue 00 is at least $\chi-1=c-1$.
@@
Figure 3: The two $3$-uniform hyperflowers $H_{3,1}^{3}$ (left) and $H_{3,2}^{3}$ (right) with $3$ petals and with $1$ and $2$ central vertices, respectively.
@@
$e\subseteq V\text{ is an edge of }\Gamma^{c}_{s_{1},\ldots,s_{k}}\iff|e|=c\text{ and }|e\cap V_{i}|\leq 1\text{ for all }i\in\{1,\ldots,k\}.$
@@
Hence, $\Gamma^{c}_{s_{1},\ldots,s_{k}}$ contains all possible egdes $e$ of cardinality $c$ such that no two vertices in $e$ belong to the same partition class $V_{i}$.
@@
These are $k(s-1)$ linearly independent eigenfunctions with eigenvalue $1$.
Lastly, for $1\leq i<j\leq k$, the functions $g_{ij}$ from Definition 3.4 are $k-1$ linearly independent eigenfunctions with eigenvalue $\frac{k-c}{k-1}$. ∎
@@ 4.1 $d$-proper colorings of $c$-uniform unoriented hypergraphs
Consider a $c$-uniform hypergraph in which the vertices of a represent the employees of a given company, while the edges represent project teams. Suppose that at most $d$ employees can be absent from a project team without overburdening the remaining members. Then, a $d$-proper coloring corresponds to a partition of the employees into groups that can have vacation at the same time.
@@
Let $d$ be an integer such that be such that $11\leq d\leq c-$ 1. Then,
@@
$=$ $1+2\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\in V_{i}\\ w\in V_{j}\end{subarray}}\bigl{|}A_{v,w}\bigr{|}}{\sum_{v\in V}\deg v}+2\frac{\sum_{i}\sum_{\begin{subarray}{c}v,w\in V_{i}\\ v\neq w\end{subarray}}\bigl{|}A_{v,w}\bigr{|}}{\sum_{v\in V}\deg v}$
$=$ $1+2\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\in V_{i}\\ w\in V_{j}\end{subarray}}\bigl{|}A_{v,w}\bigr{|}}{\frac{1}{\chi_{d}-1}\sum_{i\neq j}\sum_{v\in V_{i}\cup V_{j}}\deg v}-2(\chi_{d}-1)\frac{\sum_{i}\sum_{\begin{subarray}{c}v,w\in V_{i}\\ v\neq w\end{subarray}}\bigl{|}A_{v,w}\bigr{|}}{\sum_{v\in V}\deg v}$
@@
$2\bigl{(}\chi_{d}-1\bigr{)}\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\in V_{i}\\ w\in V_{j}\end{subarray}}\bigl{|}A_{v,w}\bigr{|}}{\sum_{i\neq j}\sum_{v\in V_{i}\cup V_{j}}\deg v}-2(\chi_{d}-1)\frac{\bigl{(}\chi_{d}-1\bigr{)}\sum_{i}\sum_{\begin{subarray}{c}v,w\in V_{i}\\ v\neq w\end{subarray}}\bigl{|}A_{v,w}\bigr{|}}{(\chi_{d}-1)\sum_{v\in V}\deg v}$
@@
=$\overset{(i)}{=}$ $\bigl{(}\chi_{d}-1\bigr{)}\bigl{(}1-\textrm{RQ}(g_{12})\bigr{)}$
[…]
@@
where, in $(ii)$, we used the fact that, for every vertex $v\in V$ and every edge $e\ni v$, there are at most $d-1$ other vertices in $e$ that have the same color as $v$.
@@
Figure 4: The $9$-uniform hyperflower $H_{3,2}^{9}$, for which the bound in Theorem 4.3 is sharp.
@@
|$\bigl{|}\{e\in E\colon v\in e\text{ and }e\cap V_{i}\neq\varnothing\}\bigr{|}=\begin{cases}\frac{(c/m-1)\deg v}{k-1},&\text{ if }v\notin V_{i},\\ \deg v,&\text{ if }v\in V_{i}.\end{cases}$
@@
$=\frac{1}{\deg v_{i}}\biggl{(}\bigl{|}\bigl{\{}(e,w_{i})\colon v_{i}\in e,w_{i}\in e\cap V_{i}\bigr{\}}\bigr{|}-\bigl{|}\bigl{\{}(e,w_{i})\colon v_{i}\in e,w_{i}\in e\cap V_{j}\bigr{\}}\bigr{|}\biggr{)}$
$=\frac{1}{\deg v_{i}}\bigl{(}\deg v_{i}\cdot m-m\cdot\bigl{|}\bigl{\{}e\colon v_{i}\in e,e\cap V_{j}\neq\varnothing\bigr{\}}\bigr{|}\bigr{)}$
@@
Analogously, we one can show that
@@
    |$\bigl{|}\bigl{\{}e\colon v_{i}\in e\text{ and }e\cap V_{j}\neq\varnothing\bigr{\}}\bigr{|}=\frac{\deg v_{i}\bigl{(}c/d-1\bigr{)}}{\chi_{d}-1}.$
@@
In this case, the corresponding eigenvalue is $(k-c)/(k-1)$.
@@
Let $\Gamma$ be a $c$-uniform oriented hypergraph with only positive signs, and let $V_{1},\ldots,V_{\chi}$ be the coloring classes of a fixed proper strong $\chi$-coloring. If Theorem 3.1 holds with equality, then for all $v\in V$ and for all $1\leq i\leq\chi$ we have that
@@ 4.2 $q$-tailored colorings
In particular, if $\Gamma$ is a $c$-uniform hypergraph in which every vertex is an input in all edges it belongs to, then any $d$-proper coloring is also a $(d-1)$-tailored coloring. This is why $q$-tailored colorings generalize $d$-proper colorings in this setting.
@@ 5 Coloring graphs $d$-improperly
Recall. A vertex coloring $c\colon V\to\{1,\ldots,k\}$ is $d$-improper if every vertex $v$ has at most $d$ neighbors with the same color as $v$. The $d$-improper coloring number, notation $\chi^{d}=\chi^{d}(G)$, is the minimal number of colors needed to color $G$ $d$-improperly.
@@
A proper vertex coloring of a graph is 00-improper. Moreover, if $G$ is a $k$-regular graph and if $d\geq 0$, then a $d/k$-tailored coloring (Definition 2.24) corresponds to a $d$-improper coloring.
@@
00 0 $=\textrm{RQ}\bigl{(}\boldsymbol{1}\bigr{)}=\frac{\sum_{v\sim w}\bigl{(}\boldsymbol{1}(v)-\boldsymbol{1}(w)\bigr{)}^{2}}{\sum_{v\in V}\deg v\cdot\boldsymbol{1}(v)^{2}}$
@@
$=1-2\bigl{(}\chi^{d}-1\bigr{)}\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\sim w\\ v\in V_{i}\\ w\in V_{j}\end{subarray}}1}{\sum_{i\neq j}\sum_{v\in V_{i}\cup V_{j}}\deg v}+2\bigl{(}\chi^{d}-1\bigr{)}\frac{\bigl{(}\chi^{d}-1\bigr{)}\sum_{i}\sum_{\begin{subarray}{c}v\sim w\\ v,w\in V_{i}\end{subarray}}1}{\bigl{(}\chi^{d}-1\bigr{)}\sum_{v\in V}\deg v}-2\chi^{d}\frac{\sum_{i}\sum_{\begin{subarray}{c}v\sim w\\ v,w\in V_{i}\end{subarray}}1}{\sum_{v\in V}\deg v}$
@@
00 0 $\geq 1-\bigl{(}\lambda_{N}-1\bigr{)}\bigl{(}\chi^{d}-1\bigr{)}-\chi^{d}\cdot\frac{d}{\overline{\deg}}.$
@@
00 0 $=Lg_{jk}(v)=\sum_{\begin{subarray}{c}w\in V_{k}\\ w\sim v\end{subarray}}\frac{1}{\deg v}-\sum_{\begin{subarray}{c}w\in V_{j}\\ w\sim v\end{subarray}}\frac{1}{\deg v}$
@@
00 0 $=d\cdot\chi^{d}\biggl{(}1-\frac{\deg v}{\overline{\deg}}\biggr{)}.$

### v2

@@ 1 Introduction
To accommodate more flexible coloring constraints, De Werra (1979) [12] introduced $I$-regular colorings: for every edge, the number of vertices of any color that it contains is bounded both above and below. In this paper, we introduce and study a specific case of $I$-regular colorings called $d$-proper colorings: these are colorings where every edge contains at most $d$ vertices of a fixed color. The concept of $d$-proper colorings generalizes strong colorings (which correspond to the case $d=1$). Moreover, for $c$-uniform hypergraphs (i.e., hypergraphs whose edges all have the same size $c$), a c1c-1-proper coloring corresponds to a weak coloring.
@@
In this paper, we shall study spectral bounds for various notions of coloring numbers, both for classical hypergraphs and oriented hypergraphs, which were introduced by Shi (1992) [35], as a generalization of classical hypergraphs in which an orientation is attached to every vertex-edge incidence. The adjacency matrix and Kirchoff Laplacian matrices for oriented hypergraphs were introduced by Reff and Rusnak (2012) [30], and the normalized Laplacian was introduced by Jost and Mulas (2019) [21]. These matrices encode information about the structure of the hypergraphs, as they do in the graph case. We refer the reader to [7, 8, 13, 32, 18, 33, 22, 28, 29, 30, 34, 31] for literature on the adjacency and Kirchhhoff Laplacian matrices, and to [3, 21, 24, 26, 25] for literature on the normalized Laplacian matrix.
@@ 2.1 The normalized Laplacian for hypergraphs
where the incidence matrix corresponds to any arbitrary orientation of the graph, such that each edge has exactly one input and one output. Equivalently, the graph Kirchhoff Laplacian can be written as K=DAK=D-A, where $D$ is the degree matrix, whose diagonal entries are the degrees of the corresponding vertices, and $A$ is the adjacency matrix. Similarly, the normalized Laplacian of a graph $G$ is the matrix
@@
$A_{i,j}\coloneqq$ # $\#\text{ edges where $v_{i}$ and $v_{j}$ are anti-oriented }$
- # $\#\text{ edges where $v_{i}$ and $v_{j}$ are co-oriented.}$
@@
[…]
@@
Note that $L$ is self-adjoint with respect to $\langle\cdot,\cdot\rangle$, and that $\langle\cdot,\cdot\rangle$ is positive definite.
@@
Let $H$ be an $N$-dimensional vector space with a positive definite scalar product $(\cdot,\cdot)$, and let $A\colon H\rightarrow H$ be a self-adjoint linear operator. Let $\mathcal{H}_{k}$ be the family of all $k$-dimensional subspaces of $H$. Then the eigenvalues $\lambda_{1}\leq\ldots\leq\lambda_{N}$ of $A$ can be obtained by
@@ 2.2 Coloring hypergraphs
In particular, if $\Gamma$ is a $c$-uniform hypergraph in which every vertex is an input in all edges it belongs to, then any $d$-proper coloring is also a (d1)(d-1)-tailored coloring. This is why $q$-tailored colorings generalize $d$-proper colorings in this setting.
@@ 3.1 Strong vertex coloring of general oriented hypergraphs
Our proof relies on Rayleigh quotients and leads to new consequences, which are detailed in Proposition 3.2 and 3.3. These results show that, if equality holds in the bound from Theorem 3.1, then there must be a specific relation between the eigenspaces corresponding to $\lambda_{1}$ and $\lambda_{N}$. To better understand the structure of such hypergraphs, one needs to know these eigenspaces. This is why, to gain further insight, in Subsection 3.2, we shall focus on $c$-uniform unoriented hypergraphs, for which the eigenspace associated with $\lambda_{N}=c$ is known to consist of constant functions.
@@
$\overset{(i)}{=}1+\bigl(l-1\bigr)\bigl(1-\textrm{RQ}(g_{12})\bigr).$
@@
$\overset{(ii)}{=}1-\bigl(k-1\bigr)\bigl(\textrm{RQ}(h_{12})-1\bigr).$
@@
    For every eigenfunction $g$ with eigenvalue $\lambda_{N}$, and for all proper $\chi$-colorings with corresponding coloring classes $V_{1},\ldots,V_{\chi}$, we have
@@ 3.2 Strong vertex coloring of unoriented $c$-uniform hypergraphs
    If $\chi=c$, then $\lambda_{1}=0$, and the eigenvalue 0 has multiplicity at least c1c-1.
@@
Since $\lambda_{1}\geq 0$, the bound must be sharp. By Proposition 3.2, this implies that the multiplicity of the eigenvalue 0 is at least $\chi-1=c-1$.
@@
Figure 3: The two 33-uniform hyperflowers $H_{3,1}^{3}$ (left) and $H_{3,2}^{3}$ (right) with $3$ petals and with $1$ and $2$ central vertices, respectively.
@@
[…]
@@
Hence, $\Gamma^{c}_{s_{1},\ldots,s_{k}}$ contains all possible edges $e$ of cardinality $c$ such that no two vertices in $e$ belong to the same partition class $V_{i}$.
@@
These are k(s1)k(s-1) linearly independent eigenfunctions with eigenvalue $1$.
Lastly, for $1\leq i<j\leq k$, the functions $g_{ij}$ from Definition 3.4 are k1k-1 linearly independent eigenfunctions with eigenvalue $\frac{k-c}{k-1}$. ∎
@@ 4.1 $d$-proper colorings of $c$-uniform unoriented hypergraphs
Consider a $c$-uniform hypergraph in which the vertices represent the employees of a given company, while the edges represent project teams. Suppose that at most $d$ employees can be absent from a project team without overburdening the remaining members. Then, a $d$-proper coloring corresponds to a partition of the employees into groups that can have vacation at the same time.
@@
Let $d$ be an integer such that $1\leq d\leq c-1$. Then,
@@
$=$ $1+2\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\in V_{i}\\ w\in V_{j}\end{subarray}}\bigl|A_{v,w}\bigr|}{\sum_{v\in V}\deg v}+2\frac{\sum_{i}\sum_{\begin{subarray}{c}v,w\in V_{i}\\ v\neq w\end{subarray}}\bigl|A_{v,w}\bigr|}{\sum_{v\in V}\deg v}.$
By rewriting the denominator of the first fraction and substituting the coefficient $2$ in front of the second fraction by $2\chi_{d}-2(\chi_{d}-1)$, we obtain
$c=$ $1+2\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\in V_{i}\\ w\in V_{j}\end{subarray}}\bigl|A_{v,w}\bigr|}{\frac{1}{\chi_{d}-1}\sum_{i\neq j}\sum_{v\in V_{i}\cup V_{j}}\deg v}-2(\chi_{d}-1)\frac{\sum_{i}\sum_{\begin{subarray}{c}v,w\in V_{i}\\ v\neq w\end{subarray}}\bigl|A_{v,w}\bigr|}{\sum_{v\in V}\deg v}$
@@
$2\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\in V_{i}\\ w\in V_{j}\end{subarray}}\bigl|A_{v,w}\bigr|}{\frac{1}{\chi_{d}-1}\sum_{i\neq j}\sum_{v\in V_{i}\cup V_{j}}\deg v}-2(\chi_{d}-1)\frac{\sum_{i}\sum_{\begin{subarray}{c}v,w\in V_{i}\\ v\neq w\end{subarray}}\bigl|A_{v,w}\bigr|}{\sum_{v\in V}\deg v}$
$=$ $2\bigl(\chi_{d}-1\bigr)\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\in V_{i}\\ w\in V_{j}\end{subarray}}\bigl|A_{v,w}\bigr|}{\sum_{i\neq j}\sum_{v\in V_{i}\cup V_{j}}\deg v}-2(\chi_{d}-1)\frac{\bigl(\chi_{d}-1\bigr)\sum_{i}\sum_{\begin{subarray}{c}v,w\in V_{i}\\ v\neq w\end{subarray}}\bigl|A_{v,w}\bigr|}{(\chi_{d}-1)\sum_{v\in V}\deg v}$
@@
$\overset{(i)}{=}$ $\bigl(\chi_{d}-1\bigr)\bigl(1-\textrm{RQ}(g_{12})\bigr)$
${\leq}$ $\bigl(\chi_{d}-1\bigr)\bigl(1-\lambda_{1}\bigr),$
@@
where, in $(ii)$, we used the fact that, for every vertex $v\in V$ and every edge $e\ni v$, there are at most d1d-1 other vertices in $e$ that have the same color as $v$.
@@
Figure 4: The 99-uniform hyperflower $H_{3,2}^{9}$, for which the bound in Theorem 4.3 is sharp.
@@
$\bigl|\{e\in E\colon v\in e\text{ and }e\cap V_{i}\neq\varnothing\}\bigr|=\begin{cases}\frac{(c/m-1)\deg v}{k-1},&\text{ if }v\notin V_{i},\\ \deg v,&\text{ if }v\in V_{i}.\end{cases}$ (19)
@@
$=\frac{1}{\deg v_{i}}\biggl(\bigl|\bigl\{(e,w_{i})\colon v_{i}\in e,w_{i}\in e\cap V_{i}\bigr\}\bigr|-\bigl|\bigl\{(e,w_{i})\colon v_{i}\in e,w_{i}\in e\cap V_{j}\bigr\}\bigr|\biggr),$
where the last equality follows from the definition of the adjacency matrix and from the observation that $v_{i}$ appears in exactly $\deg v_{i}$ ordered pairs $(e,v_{i})$, ranging over all edges $e$ such that $v_{i}\in e$.
Next, recall our assumption that for every edge $e\in E$ and for each index $i=1,\ldots,k$, one has $\lvert e\cap V_{i}\rvert\in\{0,m\}$. In particular, if $v_{i}\in e\cap V_{i}$, then the intersection $e\cap V_{i}$ must contain exactly $m$ vertices, which implies the existence of precisely m1m-1 additional vertices $w\in e\cap V_{i}$ distinct from $v_{i}$. Combining this observation with (19), we obtain
$Lg_{ij}(v_{i})$ $=\frac{1}{\deg v_{i}}\bigl(\deg v_{i}\cdot m-m\cdot\bigl|\bigl\{e\colon v_{i}\in e,e\cap V_{j}\neq\varnothing\bigr\}\bigr|\bigr)$
@@
Analogously, one can show that
@@
    $\bigl|\bigl\{e\colon v_{i}\in e\text{ and }e\cap V_{j}\neq\varnothing\bigr\}\bigr|=\frac{\deg v_{i}\bigl(c/d-1\bigr)}{\chi_{d}-1}.$
@@
In this case, the corresponding eigenvalue is (kc)/(k1)(k-c)/(k-1).
@@
Let $\Gamma$ be a $c$-uniform oriented hypergraph with only positive signs, and let $V_{1},\ldots,V_{\chi}$ be the coloring classes of a fixed proper strong $\chi$-coloring. If Theorem 3.1 holds with equality, then for all $v\in V$ and for all $1\leq i\leq\chi$, we have that
@@ 4.2 $q$-tailored colorings
In particular, if $\Gamma$ is a $c$-uniform hypergraph in which every vertex is an input in all edges it belongs to, then any $d$-proper coloring is also a (d1)(d-1)-tailored coloring. This is why $q$-tailored colorings generalize $d$-proper colorings in this setting.
@@ 5 Coloring graphs $d$-improperly
Recall. A vertex coloring $c\colon V\to\{1,\ldots,k\}$ is $d$-improper if every vertex $v$ has at most $d$ neighbors with the same color as $v$. The $d$-improper coloring number, denoted $\chi^{d}=\chi^{d}(G)$, is the minimal number of colors needed to color $G$ $d$-improperly.
@@
A proper vertex coloring of a graph is 0-improper. Moreover, if $G$ is a $k$-regular graph and if $d\geq 0$, then a $d/k$-tailored coloring (Definition 2.24) corresponds to a $d$-improper coloring.
@@
$0$ $=\textrm{RQ}\bigl(\boldsymbol{1}\bigr)=\frac{\sum_{v\sim w}\bigl(\boldsymbol{1}(v)-\boldsymbol{1}(w)\bigr)^{2}}{\sum_{v\in V}\deg v\cdot\boldsymbol{1}(v)^{2}}$
@@
By substituting the coefficient $2$ in front of the second fraction by $2\chi^{d}-2(\chi^{d}-1)$, we obtain
$0$ $=1-2\bigl(\chi^{d}-1\bigr)\frac{\sum_{i\neq j}\sum_{\begin{subarray}{c}v\sim w\\ v\in V_{i}\\ w\in V_{j}\end{subarray}}1}{\sum_{i\neq j}\sum_{v\in V_{i}\cup V_{j}}\deg v}+2\bigl(\chi^{d}-1\bigr)\frac{\bigl(\chi^{d}-1\bigr)\sum_{i}\sum_{\begin{subarray}{c}v\sim w\\ v,w\in V_{i}\end{subarray}}1}{\bigl(\chi^{d}-1\bigr)\sum_{v\in V}\deg v}-2\chi^{d}\frac{\sum_{i}\sum_{\begin{subarray}{c}v\sim w\\ v,w\in V_{i}\end{subarray}}1}{\sum_{v\in V}\deg v}$
@@
$0$ $\geq 1-\bigl(\lambda_{N}-1\bigr)\bigl(\chi^{d}-1\bigr)-\chi^{d}\cdot\frac{d}{\overline{\deg}}.$
@@
$0$ $=Lg_{jk}(v)=\sum_{\begin{subarray}{c}w\in V_{k}\\ w\sim v\end{subarray}}\frac{1}{\deg v}-\sum_{\begin{subarray}{c}w\in V_{j}\\ w\sim v\end{subarray}}\frac{1}{\deg v}$
@@
$0$ $=d\cdot\chi^{d}\biggl(1-\frac{\deg v}{\overline{\deg}}\biggr).$